# FlowPrecision: Advancing FPGA-Based Real-Time Fluid Flow Estimation with Linear Quantization

### v1

@@ Abstract
Index Terms:
Flow Estimation, Soft Sensors, In-situ AI, Quantization, FPGA-based Inference, Embedded Systems
@@ II Related Work
Accurate flow estimation using soft sensors requires high-bit quantization to ensure precision, mainly due to its regression-oriented nature. Ling et al. [6] advanced in this area by quantizing an MLP model for FPGA platforms, employing a uniform power-of-2 scaling factor in their 8-bit fixed-point quantization approach. While their approach showcased FPGA deployments’ speed and energy efficiency, it faltered in precision when compared to linear quantized models on Microcontroller Units (MCUs). The primary limitation of fixed-point quantization stems from its assumption of uniform tensor distribution, leading to notable precision loss amidst real-world data variability. In our pursuit to enhance model precision for soft sensor applications on FPGAs, we also examined the work of Jain et al. [11]. Their introduction of per-tensor scaling considerably ameliorated compatibility across diverse hardware platforms and more effectively adapted to the variability inherent in tensor distributions. Beyond refining the fixed-point quantization, linear quantization is a formidable alternative. This technique applies a zero point and scaling factor to convert floating-point tensors into integer tensors linearly. Gaining traction in prominent deep learning frameworks like TensorFlow and PyTorch, linear quantization is epitomized by the Brevitas library [10] in PyTorch, showcasing enhanced model precision through its application. tgbhu8hu86tg
@@ III Fundamentals
Quantization in numerical representation involves converting continuous real numbers from the domain $\mathbb{R}$ into discrete values within the quantized domain $\mathbb{Q}$. Consider a tensor $X$, composed of real-valued elements $x$, and its quantized counterpart $X_{q}$. As depicted in Equation 1, the scale factor $S$, a floating-point parameter, defines the relationship between $X$ and $X_{q}$, while the zero point $Z$ is a fixed-point parameter representing zero in $X$. The rounding operation approximates $x$ to the nearest integer, and the clamp function ensures $x_{q}$ remains within the bounds of a $b$-bit signed integer $(-2^{b-1},2^{b-1}-1)$. De-quantization, detailed in Equation 2, reverses this process, converting $X_{q}$ back into an approximate real-valued tensor $X^{\prime}$, using the same scale factor $S$ and zero point $Z$.
$x\mapsto x_{q}$ $=\text{clamp}\left(\text{round}(\frac{x}{S})+Z,-2^{b-1},2^{b-1}-1\right)$ (1)
@@ III-B Linear Quantization
In addressing the limitations of fixed-point quantization, our approach utilizes linear quantization with adaptive parameters, as per [9]. This method dynamically computes the scaling factors $S$ and zero points $Z$ during training, customizing them to each tensor’s distribution. Equations 3 and 4 illustrate how these parameters are computed based on the tensor’s observed minimum $\alpha$ and maximum $\beta$ values. This adaptive approach ensures a more accurate numerical representation, especially for data distributions that are not symmetric around zero, thereby enhancing model performance by minimizing quantization errors and optimizing bit precision utilization.
@@ IV-A1 QAT with Linear Quantization
Quantization-Aware Training (QAT) is a favored quantization strategy that enhances ANNs’ resilience to quantization errors. This technique simulates the quantization process during training, incorporating both quantization and de-quantization steps. The non-differentiable nature of rounding operations in backpropagation is addressed using the Straight-Through Estimator [3]. We transition from the theoretical underpinnings of QAT to its practical application in our model, identifying the primary quantization objects, as detailed in Table IV-A1. These include the inputs $X$, weights $W^{1}$, biases $B^{1}$, and activations $A_{1}$ for the hidden layer, and inputs $A_{2}$, weights $W_{2}$, biases $B_{2}$, and final outputs $Y$ for the output layer. Unlike fixed-point quantization with static scale factors, our approach dynamically calculates the quantization parameters after each training batch, aligning them with the actual data distribution to minimize quantization errors. The specific quantization parameters for each object are concisely summarized in Table IV-A1. For example, the quantization scale factor for the hidden layer’s inputs is $S_{X}$, with a corresponding zero point of $Z_{X}$. Notably, all biases have no zero points for computational simplicity. Although our implementation supports mixed-precision quantization, all quantization objects in this study are uniformly quantized at 8 bits for model consistency and simplified analysis.
TABLE I: Quantization Description Layers Quantization Objects Quantization Parameters Hidden Layer $X$ $S_{X}$, $Z_{X}$ $W^{1}$ $S_{W^{1}}$, $Z_{W^{1}}$ $B^{1}$ $S_{B^{1}}$ $A^{1}$ $S_{A^{1}}$, $Z_{A^{1}}$ ReLU $A^{1}$ $S_{A^{1}}$, $Z_{A^{1}}$ $A^{2}$ $S_{A^{1}}$, $Z_{A^{1}}$ Output Layer $A^{2}$ $S_{A^{1}}$, $Z_{A^{1}}$ $W^{2}$ $S_{W^{2}}$, $Z_{W^{2}}$ $B^{2}$ $S_{B^{2}}$ $Y$ $S_{Y}$, $Z_{Y}$
@@ IV-A2 Enhanced Integer-only Model Inference
We employ the integer-only linear layer calculation method proposed by Benoit et al. [5]. This method is applied starting from the hidden layer of our model. The computation leverages quantization parameters determined during the QAT phase, which are detailed in Equation 5. To streamline the calculation, we approximate the bias term as $S_{B^{1}}B^{1}_{q}\approx S_{X}S_{W^{1}}B^{1*}_{q}$, leading to the formulation presented in Equation 6. Notably, the term $\frac{S_{X}S_{W_{1}}}{S_{A^{1}}}$ in this equation is the sole floating-point component. To maintain our commitment to integer-only computations, we employ bit-shift operations for approximating this term, as elaborated in Equation 8. Specifically, we first convert the floating-point term into a positive integer $M_{0}$, followed by a right bit-shift operation by $n$ positions to closely approximate the original floating-point value. This integer-only computation approach is also consistently applied to the output layer, as indicated in Equation 9, ensuring a uniform integer-only processing methodology throughout our model.
@@ IV-B1 Linear Layer Optimization
Our approach to optimizing the linear layer architecture draws inspiration from the designs presented in [1]. We have adapted their VHDL templates to meet our specific needs for integer-only inference.
@@ VI-C Stability Assessment on Linear Quantization
Data Sets Hidden Neurons MSE
@@ VI-D Performance on FPGA Deployment
Our final set of experiments assessed the performance of models deployed on FPGAs, with a focus on inference time, power consumption, and energy efficiency. As detailed in Table VI, the M-Linear model consistently demonstrated a modest reduction in inference time compared to the M-Fixed model, across various hidden neuron configurations on DS1. Notably, this increased processing speed was accompanied by an elevated power consumption, often exceeding a 10% increase. This trade-off between inference speed and power consumption is attributed to the pipelined parallelization in the linear layer implementation. The resource utilization, outlined in Table VII, shows that the M-Linear model requires more Look-Up Tables (LUTs) and Block RAM (BRAM), especially as the neuron count increases. The higher use of Digital Signal Processing slices (DSPs) is necessary for the complex scaling logic, particularly evident at higher model complexities, as detailed in Algorithm 1, specifically at step 12. The uptick in resource use correlates with the model’s complexity, indicating that linear quantization enhances inference speed through more intensive use of FPGA resources. Despite the increased power consumption, the energy cost per inference of the M-Linear model is only up to 5.71% higher than the M-Fixed model. Notably, this energy consumption is still quite lower, by 56.8$\times$, compared to model inference on MCUs, as reported in [6].
@@
M-Linear 10 1.01 ($\downarrow$2.88%) 31 ($\uparrow$10.71%) 0.03 (0.0%)
@@ VII Conclusion and Future Work
Acknowledgments. The authors gratefully acknowledge the financial support provided by the Federal Ministry of Economic Affairs and Climate Protection of Germany for the RIWWER project (01MD22007C).

### v3

@@ II Related Work
Accurate flow estimation using soft sensors requires high-bit quantization to ensure precision, mainly due to its regression-oriented nature. Ling et al. [5] advanced in this area by quantizing an MLP model for FPGA platforms, employing a uniform power-of-2 scaling factor in their 8-bit fixed-point quantization approach. While their approach showcased FPGA deployments’ speed and energy efficiency, it faltered in precision when compared to linear quantized models on Microcontroller Units (MCUs). The primary limitation of fixed-point quantization stems from its assumption of uniform tensor distribution, leading to notable precision loss amidst real-world data variability. In our pursuit to enhance model precision for soft sensor applications on FPGAs, we also examined the work of Jain et al. [6]. Their introduction of per-tensor scaling considerably ameliorated compatibility across diverse hardware platforms and more effectively adapted to the variability inherent in tensor distributions. Beyond refining the fixed-point quantization, linear quantization is a formidable alternative. This technique applies a zero point and scaling factor to convert floating-point tensors into integer tensors linearly. Gaining traction in prominent deep learning frameworks like TensorFlow and PyTorch, linear quantization is epitomized by the Brevitas library [7] in PyTorch, showcasing enhanced model precision through its application.
@@ III Fundamentals
Quantization in numerical representation involves converting continuous real numbers from the domain $\mathbb{R}$ into discrete values within the quantized domain $\mathbb{Q}$. Consider a tensor $X$, composed of real-valued elements $x$, and its quantized counterpart $X_{q}$. As depicted in Equation 1, the scale factor $S$, a floating-point parameter, defines the relationship between $X$ and $X_{q}$, while the zero point $Z$ is a fixed-point parameter representing zero in $X$. The rounding operation approximates $x$ to the nearest integer, and the clamp function ensures $x_{q}$ remains within the representable range $q_{\min},q_{\max}]$ defined by the chosen quantized format, where $q_{\min}$ and $q_{\max}$ denote the lower and upper representable limits of the quantized domain. De-quantization, detailed in Equation 2, reverses this process, converting $X_{q}$ back into an approximate real-valued tensor $X^{\prime}$, using the same scale factor $S$ and zero point $Z$.
$X\mapsto X_{q}$ $=\text{clamp}\left(\text{round}(\frac{X}{S})+Z,q_{min},q_{max}\right)$ (1)
@@ III-B Linear Quantization
In addressing the limitations of fixed-point quantization, our approach utilizes linear quantization with adaptive parameters, as per [8]. This method dynamically computes the scaling factors $S$ and zero points $Z$ during training, customizing them to each tensor’s distribution. Equations 3 and 4 illustrate how these parameters are computed based on the tensor’s observed minimum $\alpha$ and maximum $\beta$ values. Here, we adopt a signed integer target with bit-width $b$, where the representable range is defined as $q_{\min}\!=\!-2^{b-1}$ and $q_{\max}\!=\!2^{b-1}-1$. This adaptive approach ensures a more accurate numerical representation, especially for data distributions that are not symmetric around zero, thereby enhancing model performance by minimizing quantization errors and optimizing bit precision utilization.
@@ IV-A1 QAT with Linear Quantization
Quantization-Aware Training (QAT) is a favored quantization strategy that enhances ANNs’ resilience to quantization errors. This technique simulates the quantization process during training, incorporating both quantization and de-quantization steps. The non-differentiable nature of rounding operations in backpropagation is addressed using the Straight-Through Estimator [9]. We transition from the theoretical underpinnings of QAT to its practical application in our model, identifying the primary quantization objects, as detailed in Table LABEL:tab:quantization_parameters. These include the inputs $X$, weights $W^{1}$, biases $B^{1}$, and activations $A_{1}$ for the hidden layer, and inputs $A_{2}$, weights $W_{2}$, biases $B_{2}$, and final outputs $Y$ for the output layer. Unlike fixed-point quantization with static scale factors, our approach dynamically calculates the quantization parameters after each training batch, aligning them with the actual data distribution to minimize quantization errors. The specific quantization parameters for each object are concisely summarized in Table LABEL:tab:quantization_parameters. For example, the quantization scale factor for the hidden layer’s inputs is $S_{X}$, with a corresponding zero point of $Z_{X}$. Notably, all biases have no zero points for computational simplicity. Although our implementation supports mixed-precision quantization, all quantization objects in this study are uniformly quantized at 8 bits for model consistency and simplified analysis.
TABLE I: Quantization Description
Layers
Quantization
Objects
Quantization
Parameters
Hidden Layer $X$ $S_{X}$, $Z_{X}$
$W^{1}$ $S_{W^{1}}$, $Z_{W^{1}}$
$B^{1}$ $S_{B^{1}}$
$A^{1}$ $S_{A^{1}}$, $Z_{A^{1}}$
ReLU $A^{1}$ $S_{A^{1}}$, $Z_{A^{1}}$
$A^{2}$ $S_{A^{1}}$, $Z_{A^{1}}$
Output Layer $A^{2}$ $S_{A^{1}}$, $Z_{A^{1}}$
$W^{2}$ $S_{W^{2}}$, $Z_{W^{2}}$
$B^{2}$ $S_{B^{2}}$
$Y$ $S_{Y}$, $Z_{Y}$
@@ IV-A2 Enhanced Integer-only Model Inference
We employ the integer-only linear layer calculation method proposed by Benoit et al. [10]. This method is applied starting from the hidden layer of our model. The computation leverages quantization parameters determined during the QAT phase, which are detailed in Equation 5. To streamline the calculation, we approximate the bias term $S_{B^{1}}B^{1}_{q}$ as $S_{X}S_{W^{1}}B^{1*}_{q}$, leading to the formulation presented in Equation 6. Notably, the term $\frac{S_{X}S_{W_{1}}}{S_{A^{1}}}$ in this equation is the sole floating-point component. To maintain our commitment to integer-only computations, we employ bit-shift operations for approximating this term, as elaborated in Equation 8. Specifically, we first convert the floating-point term into a positive integer $M_{0}$, followed by a right bit-shift operation by $n$ positions to closely approximate the original floating-point value. This integer-only computation approach is also consistently applied to the output layer, as indicated in Equation 9, ensuring a uniform integer-only processing methodology throughout our model.
@@ IV-B1 Linear Layer Optimization
Our approach to optimizing the linear layer draws inspiration from the designs presented in [11]. We have adapted their VHDL templates to meet our specific needs for integer-only inference.
@@ VI-C Stability Assessment on Linear Quantization
Data Sets
Hidden
Neurons
MSE
@@ VI-D Performance on FPGA Deployment
[…]
@@
M-Linear
10 1.01 ($\downarrow$2.88%) 31 ($\uparrow$10.71%) 0.03 (0.0%)
@@ VII Conclusion and Future Work
Acknowledgments. The authors gratefully acknowledge the financial support provided by the Federal Ministry for Economic Affairs and Climate Action of Germany for the RIWWER project (01MD22007C).